OpenStudy (amtran_bus):

At a snack bar, juice costs twice as much as bacon, and an order of 3 juices and 2 bacons costs $20, how much does each juice cost? SET UP AS A SERIES OF EQUTIONS

OpenStudy (amtran_bus):

@Jhannybean you are faithful

OpenStudy (anonymous):

j=2b letj=juice and b=bacon 3j+2b=20 substitue for j 3(2b)+2b=20 6b+2b=20 8b=20 divide by 8 b=2.50 now substitue for b j=2(2.5) j=5 now substitue for j 3(3)+2b=20 9+2b=20 2b=11 divide by 2 b=5.5 so bacon is 5.50 and juice is 5 understand?

OpenStudy (amtran_bus):

Thank you!!

OpenStudy (anonymous):

welcome